A bag contains 9 red marbles, 5 white marbles, and 7 blue marbles. You draw 5 marbles out at random, without replacement. What is the probability that exactly 2 are red
Answer #1

Solution:

Given in the question

there are 9 red marbles, 5 white marbles and 7 blue marbles so there is 21 marbles

P = the number of wanted outcomes/ no. of possible outcomes

P = 2 red balls and 3 other than red balls/ No. of possible outcomes = (9*8*12*11*10)/(21*20*19*18*17) = 95040/2441880 = 0.0389